### v1

@@ 5 Threats to Validity
Another potential source of threat could be the replication of existing work. The original replication package was unavailable; hence, we used the publicly available version of BugLocator, BLUiR [49], and DNNLOC [50]. For BLIA, we reused the author’s replication package [23]. We validated our implementation of the existing methods using their original dataset and achieved comparable results (e.g., with differences $\approx$ 2.00%–3.00% using MAP).
Threats to conclusion validity. The observations from our study and the conclusions we drew from them could be a source of threat to conclusion validity [73]. In this research, we answer three research questions using two different datasets and re-implement four existing techniques. We use appropriate statistical tests (e.g., t-test) and report the test details (e.g., p-value, Cohen’s D) to conclude. Thus, such threats might also be mitigated.
@@ 6.3 Deep learning-based bug localization
To address the above gap, in this empirical study, we replicated four existing techniques [11, 12, 23] to locate bugs in deep learning systems. Unlike Kim et al. [22], our study extends beyond bug localization from deep learning systems. Our study evaluates existing bug localization techniques, categorizes deep-learning bugs, analyzes their prevalence and challenges, and assesses each technique’s effectiveness for different bug types. We also conduct extensive manual analysis and explain they are difficult to localize (e.g., extrinsic factors, multifaceted dependencies), which makes our work novel.
@@ 7 Conclusion
Identifying the location of a bug within a software system (a.k.a. bug localization) is crucial to correct any bug. In recent years, bug localization techniques have received considerable attention in the context of traditional software systems. However, they might not be sufficient for deep learning systems as deep learning bugs pose a greater challenge due to their multifaceted dependencies. However, the potential of existing approaches for localizing bugs in deep learning systems is not well understood to date. In this work, we replicated four existing bug localization approaches and found that they show poor performance in localizing bugs from deep-learning systems. Secondly, through an in-depth analysis, we found that localizing certain categories of bugs (e.g., training bugs & GPU bugs) is more difficult than other bugs in deep learning systems. Finally, we investigate and find that deep learning bugs are more likely to be extrinsic, i.e., connected to non-code artifacts (e.g., training data). Our research thus offers empirical evidence and actionable insights for deep learning software bugs, advancing automated software debugging research. Future work can focus on developing a new framework for automated software debugging based on the insights from this empirical study.
Data Availability Statement (DAS)
All the data generated or analyzed during this study are available in the GitHub Repository to help reproduce our results [46].
Conflict of Interest
The authors declare that they have no conflict of interest.

### v2

@@ 5 Threats to Validity
Another potential source of threat could be the replication of existing work. The original replication package was unavailable; hence, we used the publicly available version of BugLocator, BLUiR [54], and DNNLOC [65]. For BLIA and bjXnet, we reused the author’s replication package [23, 30]. We validated our implementation of the existing methods using their original dataset and achieved comparable results (e.g., with differences $\approx$ 2.00%–3.00% using MAP).
Threats to conclusion validity. The observations from our study and the conclusions we drew from them could be a source of threat to conclusion validity [96]. In this research, we answer three research questions using two different datasets and re-implement five existing techniques. We use appropriate statistical tests (e.g., t-test) and report the test details (e.g., p-value, Cohen’s D) to conclude. Thus, such threats might also be mitigated.
@@ 6.1 Software bug
6.2 Spectrum-based bug localization
Spectrum-based bug localization techniques have been widely studied for bug localization in traditional software systems. These techniques analyze program traces from passing and failing tests to measure how frequently each program element is executed in both cases [78]. Based on these frequencies, spectrum-based techniques rank program elements by their suspiciousness, with the aim of prioritizing fault-prone locations. For example, Tarantula [55] is a widely known spectrum-based localization technique that calculates suspiciousness by measuring the ratio of failed to total executions of a particular element, highlighting the ones more frequently associated with failures.
However, capturing execution traces for spectrum-based approaches is not as straightforward in deep learning systems. One of the main reasons for this is that not all bugs in DL systems result in crashes or stack traces [36]. Instead, many DL system bugs manifest as silent bugs [57], leading to incorrect but seemingly valid behavior, making it difficult to generate program traces that are traditionally used in spectrum-based techniques. Moreover, the non-deterministic and data-driven nature of DL models further complicates the application of spectrum-based bug localization techniques. For instance, DeepFault [56] attempts to apply spectrum-based techniques to Deep Neural Networks (DNNs) by focusing on neuron activation spectra to identify suspicious neurons. While this approach is promising, it fails to handle more complex bugs, such as silent bugs or those caused by external factors, including GPU-related issues.
Another major challenge of applying spectrum-based bug localization techniques to deep learning systems is the resource-intensive nature of these systems. DL models are computationally expensive to train and execute, requiring substantial hardware resources, such as GPUs and large memory capacities. The overhead of collecting and processing execution traces for spectrum-based approaches can, therefore, become prohibitive, particularly in large-scale DL projects. Given these limitations, spectrum-based bug localization techniques are not well-suited for our study, which focuses on localizing bugs in DL systems.
@@ 6.3 Information Retrieval-based bug localization
To the best of our knowledge, no IR-based techniques have been developed exclusively for localizing bugs in DL systems. Existing approaches for bug detection in DL systems, such as DeepLocalize [58], DeepDiagnosis [20], and DeepFD [21], focus on specific types of DL bugs (e.g., model bugs, training bugs) and rely on dynamic analysis or heuristic-based methods. While these approaches are valuable, they are limited in terms of scope, scalability, and generalizability. They cannot be applied to non-DL bugs. For example, DeepLocalize is tightly coupled with the Keras library and exhibits low accuracy in bug detection. DeepDiagnosis depends on hard-coded rules, restricting its adaptability across different contexts. Moreover, these techniques are highly specific to DL models and cannot be adapted to traditional software bugs, making direct comparisons infeasible. On the other hand, IR-based techniques like BugLocator, BLUiR, and BLIA are more general-purpose and can be applied to both DL and non-DL systems.
By including traditional IR-based and recent deep learning-based baselines in our study, we attempt to identify the strengths and limitations of these techniques when applied to DL systems.
@@ 6.4 Deep learning-based bug localization
Additionally, Cao et al. [59] characterize the performance problems in deep learning (DL) systems and identify key stages where these problems are introduced and exposed. They highlight the unique challenges of detecting and localizing bugs in DL systems, particularly due to dependencies on GPUs and external libraries like TensorFlow and Keras. However, their focus is limited to performance problems related to hardware dependencies, overlooking broader system components such as DNN model architecture and training dynamics. For example, they do not address how issues like vanishing gradients during model training or incorrect hyperparameter configurations (e.g., learning rate or batch size) can also lead to performance degradation in deep learning systems. Therefore, empirical research like ours, which focuses on the challenges posed by deep learning bugs in the software debugging process (specifically in bug localization), is essential.
To address the above gap, in this empirical study, we replicated five existing techniques [11, 12, 23, 17, 30] to locate bugs in deep learning systems. Unlike Kim et al. [22], our study extends beyond bug localization from deep learning systems. Our study evaluates existing bug localization techniques, categorizes deep-learning bugs, analyzes their prevalence and challenges, and assesses each technique’s effectiveness for different bug types. We also conduct extensive manual analysis and explain they are difficult to localize (e.g., extrinsic factors, multifaceted dependencies), which makes our work novel.
@@ 7 Conclusion
Identifying the location of a bug within a software system (a.k.a. bug localization) is crucial to correct any bug. In recent years, bug localization techniques have received considerable attention in the context of traditional software systems. However, they might not be sufficient for deep learning systems as deep learning bugs pose a greater challenge due to their multifaceted dependencies. However, the potential of existing approaches for localizing bugs in deep learning systems is not well understood to date. In this work, we replicated five existing bug localization approaches and found that they show poor performance in localizing bugs from deep-learning systems. Secondly, through an in-depth analysis, we found that localizing certain categories of bugs (e.g., training bugs & GPU bugs) is more difficult than other bugs in deep learning systems. Finally, we investigate and find that deep learning bugs are more likely to be extrinsic, i.e., connected to non-code artifacts (e.g., training data). Our research thus offers empirical evidence and actionable insights for deep learning software bugs, advancing automated software debugging research. Future work can focus on developing a new framework for automated software debugging based on the insights from this empirical study.
Statement and Declarations
Funding
This work was supported by the Natural Sciences and Engineering Research Council of Canada (Discovery Grant RGPIN‑03236).
Conflicts of Interests
The authors declare they have no conflict of interest.
Ethical Approval
Not applicable.
Informed Consent
Not applicable.
Author Contributions
Conceptualization: Sigma Jahan; Methodology: Sigma Jahan; Formal analysis and investigation: Sigma Jahan; Writing—original draft preparation: Sigma Jahan; Writing—review & editing: Sigma Jahan; Resources: Sigma Jahan, Mohammad Masudur Rahman; Supervision: Mohammad Masudur Rahman; Data annotation: Sigma Jahan, Mehil B Shah; Replication of baseline technique: Sigma Jahan, Mehil B Shah.
Data Availability Statement
The dataset and supplementary materials are publicly available at https://doi.org/10.5281/zenodo.15091574.
Appendix A Appendix
Figure 12: Code snippet of the ground truth for the model bug [40] (https://bit.ly/3XJCdqq)
Figure 13: Code snippet of the incorrect source file for the model bug retrieved by BugLocator (https://bit.ly/43klnj9)
Figure 14: Code snippet of the incorrect source file for the model bug retrieved by BLUiR (https://bit.ly/3PQo88U)
Figure 15: Code snippet of the ground truth for training bug [41] (https://bit.ly/3NKfRAH)
Figure 16: Code snippet of the incorrect source file for the training bug retrieved by BugLocator (https://bit.ly/3O5yEIa)
Figure 17: Code snippet of the ground truth for the tensor bug [42] (https://bit.ly/3XLe82y)
Figure 18: Code snippet of the incorrect source file for the tensor bug retrieved by BugLocator (https://bit.ly/3Dr5LA9)
Figure 19: Code snippet of the ground truth for the API bug [43] (https://bit.ly/3JT1JE8)
Figure 20: Code snippet of the ground truth for the GPU bug [44] (https://bit.ly/44llZXc)